# Graph Limits via Quotients

### v1

@@ 4.4 Missing Proofs from Section 4
By construction, we have $\mathsf{M}^{\prime}\overset{d}{=}\mathsf{M}$ and $\mathsf{G}^{(n)}\overset{d}{=}\mathsf{G}^{(n)}[\mathsf{M}]$. Next, because $\mathsf{M}$ is exchangeable, we claim that $t_{1},\ldots,t_{N}\overset{iid}{\sim}\mathrm{Unif}([0,1])$. Indeed, $(x_{i},y_{i})\overset{d}{=}(\sigma(x_{i}),\sigma(y_{i}))$ for all $\sigma\in S_{[0,1]}$, hence $(x_{i},y_{i})\sim\theta\lambda^{2}+(1-\theta)\lambda_{D}$. Therefore, for each $i$ either $x_{i}=y_{i}\sim\mathrm{Unif}([0,1])$ or $x_{i},y_{i}\overset{iid}{\sim}\lambda^{2}$. Combining this with the fact that the $(x_{i},y_{i})$ are independent for different $i$, we obtain the claim. Therefore, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}\overset{d}{=}\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}=\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}$, the latter being an empirical measure drawn from $\mu\sim\mathsf{M}$. Finally, we have $\mathsf{M}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}$ and $\mathsf{M}_{\mathsf{G}^{(n)}}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}_{\mathsf{G}^{(n)}}$ as can be seen from (16), as $\mathsf{G}^{(n)}$ only depends on the multiplicities of $\{(t_{i},t_{j})\}$ rather than their locations, and the multiplicities and locations of a random exchangeable measure are independent of each other [kallenberg1990exchangeable, Thm. 2]. ∎
@@
Fix any $\mu\in\mathcal{P}([0,1]^{2})$. By the proof of [wainwright2019high, Thm. 4.10], we have
@@
where $R_{n}(\mathcal{F})=\underset{\begin{subarray}{c}X_{i}\overset{iid}{\sim}\mu\\ \varepsilon_{i}\overset{iid}{\sim}\mathrm{Unif}(\{\pm 1\})\end{subarray}}{\mathbb{E}}\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|$ is the Rademacher complexity of a class of functions $\mathcal{F}$ and we set $\mathcal{F}_{\mathcal{AAR}}=\{\mathbbm{1}_{R}:R\in\mathcal{AAR}\}$. We further have by [wainwright2019high, Ex. 5.24] that
@@
where $N(t;\mathcal{F};\|\cdot\|_{\mu_{n}})$ is the $t$-covering number of $\mathcal{F}$ in the (random) metric $\|f-g\|_{\mu_{n}}=\sqrt{\mathbb{E}_{\mu_{n}}(f-g)^{2}}$ defined by the empirical measure $\mu_{n}$. If $\mathcal{F}$ is class of $\{0,1\}$-valued functions, then $\mathrm{diam}(\mathcal{F})\leq 1$. Finally, we have by [HAUSSLER1995217, Cor. 1] that
@@ 5.1 Equipartition Consistency
We begin by arguing that $(\mathsf{G}_{k}[\mathsf{M}])$ is equipartition-consistent. Note that $d_{k,nk}\circ F_{nk,m}\overset{d}{=}F_{k,m}$ for any $k,n,m\in\mathbb{N}$ by [levin2025deFin, Lemma 4.13], since each $i\in[m]$ is mapped independently by $F_{nk,m}$ to one of the fibers of $d_{k,nk}$ with the same probability. Therefore, for any $G\in\Delta^{n\times n}$ the sequence of distributions $(\mathrm{Law}(\rho(F_{k,n})G_{n}))_{k}$ is indeed an equipartition-consistent model. If $(G_{m})$ is a sequence of finite graphs converging to a grapheur $\mathsf{M}$, then $\rho(F_{k,m})G_{m}\to\mathsf{G}_{k}[\mathsf{M}]$ and $\rho(F_{nk,m})G_{m}\to\mathsf{G}_{nk}[\mathsf{M}]$ weakly, hence
@@
Fix a grapheur $\mathsf{M}\in\mathcal{M}$ and consider the associated equipartition-consistent random graph model $(\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]))_{k}$.
@@
    Draw $\mu\sim\mathsf{M}$ and let $\mathsf{G}_{k}=(\mu(I_{i}^{(n)}\times I_{j}^{(n)}))_{i,j\in[k]}$ for each $k\in\mathbb{N}$. Then $G_{k}\sim\mu_{k}$ for all $k\in\mathbb{N}$, and $(G_{k})$ is quotient-convergent to $\mathsf{M}$ almost surely.
@@
    Suppose $\mathsf{M}=\delta_{(T_{1},T_{2})}$ is the grapheur associated to a single edge. Draw $G_{k}\sim\mu_{k}$ independently of each other. Then $\limsup_{k}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ almost surely, so $(G_{k})$ does not converge to $\mathsf{M}$.
@@
For the first claim, fix any $\epsilon>0$ and let $N\in\mathbb{N}$ be sufficiently large so that $\sum_{i>N\textrm{ or }j>N}E_{i,j}+\sum_{i>N}(\sigma_{i}+\varsigma_{i})\leq\epsilon$. Drawing $\mu\sim\mathsf{M}$ amounts to sampling locations $T_{i}\overset{iid}{\sim}\mathrm{Unif}([0,1])$ in (5), and these samples are distinct almost surely. Therefore, there exists $K\geq N\in\mathbb{N}$ such that $T_{1},\ldots,T_{N}$ lie in distinct intervals $(I_{i}^{(k)})_{i\in[k]}$ for all $k\geq K$. Therefore, for all $k\geq K$ we can find a permutation $\pi_{k}\in\mathsf{S}_{k}$ such that $\pi_{k}G_{k}=\rho(\iota_{k,N})(E_{i,j}+\sigma_{i}/k+\varsigma_{j}/k)_{i,j\in[N]}+N_{k}+\theta\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top}+\vartheta\frac{1}{k}I_{k}$ where $\|N_{k}\|_{1}\leq\epsilon$. Using the coupling $(\mathsf{M}_{G_{k}}^{\prime},\mathsf{M}^{\prime})$ obtained by drawing random locations $(T_{i})$, setting $\mathsf{M}^{\prime}$ to (5), and setting $\mathsf{M}^{\prime}_{G_{k}}=\sum_{i,j=1}^{N}(\pi_{k}G_{k})_{i,j}\delta_{(T_{i},T_{j})}$, we conclude that $\limsup_{k\to\infty}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\leq\epsilon$. Since $\epsilon>0$ was arbitrary, we obtain the claimed convergence.
For the second claim, note that with probability $1/k$ we have $T_{1},T_{2}\in I_{i}^{(k)}$ for the same $i\in[k]$, in which case $G_{k}=\mathrm{diag}(0,\ldots,1,\ldots,0)$ and $W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ as can be seen by setting $S=[0,1/2]$ and $T=[1/2,1]$ in (7). Since these events are independent and the sum of their probabilities $\sum_{k}k^{-1}$ diverges, Borel–Cantelli implies that $\limsup_{k}W_{\square}(\mathsf{M}_{G_{k}},\mathsf{M})\geq 1/4$ almost surely. ∎
In words, the quotients must be sampled in a manner that is related across different sizes for the samples to converge to the underlying grapheur; if we instead sample random quotients independently for each size, the resulting sequence almost surely does not converge to the underlying grapheur. In contrast, if we sample growing-sized random graphs from a graphon independently for different sizes, then the resulting sequence does converge back to the graphon almost surely as a consequence of Szemerédi’s regularity lemma [lovasz2012large, Lemma 11.8]. Proposition 5.1(2) thus implies that the analog of Theorem 1.7 fails for random quotients. More precisely, for any rate of convergence $(R(k)\in\mathbb{R}_{\geq 0})_{n\in\mathbb{N}}$ with $R(k)\to 0$, we have $\sum_{k\in\mathbb{N}}\mathbb{P}[W_{\square}(\mathsf{M},\mathsf{G}_{k}[\mathsf{M}])>R(k)]=\infty$ by Borel–Cantelli. We leave the precise analysis of convergence rates of random quotients for future work.
@@
Our goal in this section is to prove Theorem 1.9, showing that every equipartition-consistent random graph model arises from a unique random exchangeable measure. We deduce the existence of the desired random measure from results on inverse limits of random histograms [orbanz2011projective], and proving that this random measure is exchangeable proceeds via standard approximation arguments. To apply the results of [orbanz2011projective], we require the following lemma on the sequence of means of an equipartition-consistent model.
@@
Combining the above lemma with [orbanz2011projective], we are ready to prove Theorem 1.9.
@@
Uniqueness is clear since squares of the form $I_{i}^{(k)}\times I_{j}^{(k)}$ generate the Borel $\sigma$-algebra on $0,1]^{2}$. To prove existence we appeal to [orbanz2011projective, Thm. 1.1], which states that there exists a random probability measure $\mathsf{M}$ on $0,1]^{2}$ (not guaranteed to be exchangeable) satisfying $\mathsf{M}(I_{i}^{(k)}\times I_{j}^{(k)})\sim\mu_{k}$ for all $k$ if there is a determinisitic measure $\mu\in\mathcal{P}([0,1]^{2})$ satisfying $(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}=\mathbb{E}_{X\sim\mu_{k}}X$ for all $k\in\mathbb{N}$. Such a deterministic measure exists by Lemma 5.2, since
@@
We proceed to prove that the above $\mathsf{M}$ is exchangeable. To do so, it suffices to prove that for any continuous function $f\colon[0,1)^{2}\to\mathbb{R}$ and any measure-preserving bijection $\sigma$, we have $\mathbb{E}_{\mathsf{M}}f\overset{d}{=}\mathbb{E}_{\mathsf{M}}f\circ(\sigma,\sigma)$. By [kallenberg2005probabilistic, Prop. 9.1], it further suffices to prove this equality for $\sigma\in\mathsf{S}_{m}$, a permutation of the $m$ intervals $I_{1}^{(m)},\ldots,I_{m}^{(m)}$ acting by $\sigma((i-1)/m+x)=(\sigma(i)-1)/m+x$ for $x\in[0,1/m)$ for $i\in[m]$.
@@ 6 Conclusions and Future Directions
We have introduced a new notion of limits of growing graphs based on convergence of their fixed-size random quotients. We showed that this notion of convergence can be equivalently characterized in terms of convergence of graph parameters such as homomorphism numbers and quotient densities. We characterized limits of quotient-convergent graph sequences via grapheurs, which are certain random exchangeable measures on $0,1]^{2}$. The limiting grapheur of a quotient-convergent sequence of graphs describes the asymptotic distribution of edge weight in the graphs. We then presented another equivalent view of quotient convergence via edge sampling, allowing us to prove a dual analog of Szemerédi’s regularity lemma and to test properties of arbitrarily-large graphs by sampling edges from them. The number of edges needed is indpendent of the size of the graph, only depending on the desired accuracy and the Lipschitz constant of the graph parameter in question. Finally, we studied equipartition-consistent random graph models, showing that any such model corresponds to the sequence of random quotients associated to a mixture of grapheurs. We conclude by mentioning several questions suggested by our work.
@@
  2. item(Rates for random quotients)
    Is there a Szemerédi-type regularity lemma for random quotients, i.e., does $\mathbb{E}_{\mathsf{G}_{k}[\mathsf{M}]}W_{\square}(\mathsf{M}_{\mathsf{G}_{k}[\mathsf{M}]},\mathsf{M})$ converge to zero at a universal rate independent of $\mathsf{M}$? If not, how does this rate depend on $\mathsf{M}$?

### v2

@@ 4.4 Missing Proofs from Section 4
By construction, we have $\mathsf{M}^{\prime}\overset{d}{=}\mathsf{M}$ and $\mathsf{G}^{(n)}\overset{d}{=}\mathsf{G}^{(n)}[\mathsf{M}]$. Next, because $\mathsf{M}$ is exchangeable, we claim that $t_{1},\ldots,t_{N}\overset{iid}{\sim}\mathrm{Unif}([0,1])$. Indeed, $(x_{i},y_{i})\overset{d}{=}(\sigma(x_{i}),\sigma(y_{i}))$ for all $\sigma\in S_{[0,1]}$, hence $(x_{i},y_{i})\sim\theta\lambda^{2}+(1-\theta)\lambda_{D}$. Therefore, for each $i$ either $x_{i}=y_{i}\sim\mathrm{Unif}([0,1])$ or $x_{i},y_{i}\overset{iid}{\sim}\lambda^{2}$. Combining this with the fact that the $(x_{i},y_{i})$ are independent for different $i$, we obtain the claim. Therefore, $\mathsf{M}_{\mathsf{G}^{(n)}[\mathsf{M}]}\overset{d}{=}\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}=\mathsf{M}^{\prime}_{\mathsf{G}^{(n)}}$, the latter being an empirical measure drawn from $\mu\sim\mathsf{M}$. Finally, we have $\mathsf{M}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}$ and $\mathsf{M}_{\mathsf{G}^{(n)}}^{\prime}|\mathsf{G}^{(n)}\overset{d}{=}\mathsf{M}_{\mathsf{G}^{(n)}}$ as can be seen from (17), as $\mathsf{G}^{(n)}$ only depends on the multiplicities of $\{(t_{i},t_{j})\}$ rather than their locations, and the multiplicities and locations of a random exchangeable measure are independent of each other [28, Thm. 2]. ∎
@@
Fix any $\mu\in\mathcal{P}([0,1]^{2})$. By the proof of [47, Thm. 4.10], we have
@@
where $R_{n}(\mathcal{F})=\underset{\begin{subarray}{c}X_{i}\overset{iid}{\sim}\mu\\ \varepsilon_{i}\overset{iid}{\sim}\mathrm{Unif}(\{\pm 1\})\end{subarray}}{\mathbb{E}}\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(X_{i})\right|$ is the Rademacher complexity of a class of functions $\mathcal{F}$ and we set $\mathcal{F}_{\mathcal{AAR}}=\{\mathbbm{1}_{R}:R\in\mathcal{AAR}\}$. We further have by [47, Ex. 5.24] that
@@
where $N(t;\mathcal{F};\|\cdot\|_{\mu_{n}})$ is the $t$-covering number of $\mathcal{F}$ in the (random) metric $\|f-g\|_{\mu_{n}}=\sqrt{\mathbb{E}_{\mu_{n}}(f-g)^{2}}$ defined by the empirical measure $\mu_{n}$. If $\mathcal{F}$ is class of $\{0,1\}$-valued functions, then $\mathrm{diam}(\mathcal{F})\leq 1$. Finally, we have by [20, Cor. 1] that
@@ 5.1 Equipartition Consistency
We begin by arguing that $(\mathsf{G}_{k}[\mathsf{M}])$ is equipartition-consistent. Note that $d_{k,nk}\circ F_{nk,m}\overset{d}{=}F_{k,m}$ for any $k,n,m\in\mathbb{N}$ by [31, Lemma 4.13], since each $i\in[m]$ is mapped independently by $F_{nk,m}$ to one of the fibers of $d_{k,nk}$ with the same probability. Therefore, for any $G\in\Delta^{n\times n}$ the sequence of distributions $(\mathrm{Law}(\rho(F_{k,n})G_{n}))_{k}$ is indeed an equipartition-consistent model. If $(G_{m})$ is a sequence of finite graphs converging to a grapheur $\mathsf{M}$, then $\rho(F_{k,m})G_{m}\to\mathsf{G}_{k}[\mathsf{M}]$ and $\rho(F_{nk,m})G_{m}\to\mathsf{G}_{nk}[\mathsf{M}]$ weakly, hence
@@
For any grapheur $\mathsf{M}\in\mathcal{M}$, denote by $(\mu_{k}=\mathrm{Law}(\mathsf{G}_{k}[\mathsf{M}]))_{k}$ its associated equipartition-consistent random graph model.
@@
    Fix a grapheur $\mathsf{M}\in\mathcal{M}$, draw $\mu\sim\mathsf{M}$, and let $\mathsf{Q}_{k}=(\mu(I_{i}^{(n)}\times I_{j}^{(n)}))_{i,j\in[k]}$ for each $k\in\mathbb{N}$. Then $\mathrm{Law}(\mathsf{Q}_{k})=\mu_{k}$ for all $k\in\mathbb{N}$, and $(\mathsf{Q}_{k})$ is quotient-convergent to $\mathsf{M}$ almost surely.
@@
    Suppose $\mathsf{M}=\delta_{(T_{1},T_{2})}$ is the grapheur associated to a single edge. Draw $\mathsf{R}_{k}\sim\mu_{k}$ independently for different $k$. Then $\limsup_{k}W_{\square}(\mathsf{M}_{\mathsf{R}_{k}},\mathsf{M})\geq 1/4$ almost surely, so $(\mathsf{R}_{k})$ does not converge to $\mathsf{M}$.
@@
For the first claim, fix any $\epsilon>0$ and let $N\in\mathbb{N}$ be sufficiently large so that $\sum_{i>N\textrm{ or }j>N}E_{i,j}+\sum_{i>N}(\sigma_{i}+\varsigma_{i})\leq\epsilon$. Drawing $\mu\sim\mathsf{M}$ amounts to sampling locations $T_{i}\overset{iid}{\sim}\mathrm{Unif}([0,1])$ in (5), and these samples are distinct almost surely. Therefore, there exists $K\geq N\in\mathbb{N}$ such that $T_{1},\ldots,T_{N}$ lie in distinct intervals $(I_{i}^{(k)})_{i\in[k]}$ for all $k\geq K$. Therefore, for all $k\geq K$ we can find a permutation $\pi_{k}\in\mathsf{S}_{k}$ such that $\pi_{k}\mathsf{Q}_{k}=\rho(\iota_{k,N})(E_{i,j}+\sigma_{i}/k+\varsigma_{j}/k)_{i,j\in[N]}+N_{k}+\theta\frac{1}{k^{2}}\mathbbm{1}_{k}\mathbbm{1}_{k}^{\top}+\vartheta\frac{1}{k}I_{k}$ where $\|N_{k}\|_{1}\leq\epsilon$. Using the coupling $(\mathsf{M}_{\mathsf{Q}_{k}}^{\prime},\mathsf{M}^{\prime})$ obtained by drawing random locations $(T_{i})$, setting $\mathsf{M}^{\prime}$ to (5), and setting $\mathsf{M}^{\prime}_{\mathsf{Q}_{k}}=\sum_{i,j=1}^{N}(\pi_{k}\mathsf{Q}_{k})_{i,j}\delta_{(T_{i},T_{j})}$, we conclude that $\limsup_{k\to\infty}W_{\square}(\mathsf{M}_{\mathsf{Q}_{k}},\mathsf{M})\leq\epsilon$. Since $\epsilon>0$ was arbitrary, we obtain the claimed convergence.
For the second claim, note that with probability $1/k$ we have $T_{1},T_{2}\in I_{i}^{(k)}$ for the same $i\in[k]$, in which case $\mathsf{R}_{k}=\mathrm{diag}(0,\ldots,1,\ldots,0)$ and $W_{\square}(\mathsf{M}_{\mathsf{R}_{k}},\mathsf{M})\geq 1/4$ as can be seen by setting $S=[0,1/2]$ and $T=[1/2,1]$ in (7). Since these events are independent and the sum of their probabilities $\sum_{k}k^{-1}$ diverges, Borel–Cantelli implies that $\limsup_{k}W_{\square}(\mathsf{M}_{\mathsf{R}_{k}},\mathsf{M})\geq 1/4$ almost surely. ∎
In words, the quotients must be sampled in a manner that is related across different sizes for the samples to converge to the underlying grapheur; if we instead sample random quotients independently for each size, the resulting sequence almost surely does not converge to the underlying grapheur. In contrast, if we sample growing-sized random graphs from a graphon independently for different sizes, then the resulting sequence does converge back to the graphon almost surely as a consequence of Szemerédi’s regularity lemma [34, Lemma 11.8]. Proposition 5.1(2) thus implies that the analog of Theorem 1.7 fails for random quotients. More precisely, for any rate of convergence $(R(k)\in\mathbb{R}_{\geq 0})_{n\in\mathbb{N}}$ with $R(k)\to 0$, we have $\sum_{k\in\mathbb{N}}\mathbb{P}[W_{\square}(\mathsf{M},\mathsf{G}_{k}[\mathsf{M}])>R(k)]=\infty$ by Borel–Cantelli. We leave the precise analysis of convergence rates of random quotients for future work.
@@
Our goal in this section is to prove Theorem 1.9, showing that every equipartition-consistent random graph model arises from a unique random exchangeable measure. We deduce the existence of the desired random measure from results on inverse limits of random histograms [38], and proving that this random measure is exchangeable proceeds via standard approximation arguments. To apply the results of [38], we require the following lemma on the sequence of means of an equipartition-consistent model.
@@
Combining the above lemma with [38], we are ready to prove Theorem 1.9.
@@
Uniqueness is clear since squares of the form $I_{i}^{(k)}\times I_{j}^{(k)}$ generate the Borel $\sigma$-algebra on $0,1]^{2}$. To prove existence we appeal to [38, Thm. 1.1], which states that there exists a random probability measure $\mathsf{M}$ on $0,1]^{2}$ (not guaranteed to be exchangeable) satisfying $\mathsf{M}(I_{i}^{(k)}\times I_{j}^{(k)})\sim\mu_{k}$ for all $k$ if there is a determinisitic measure $\mu\in\mathcal{P}([0,1]^{2})$ satisfying $(\mu(I_{i}^{(k)}\times I_{j}^{(k)}))_{i,j\in[k]}=\mathbb{E}_{X\sim\mu_{k}}X$ for all $k\in\mathbb{N}$. Such a deterministic measure exists by Lemma 5.2, since
@@
We proceed to prove that the above $\mathsf{M}$ is exchangeable. To do so, it suffices to prove that for any continuous function $f\colon[0,1)^{2}\to\mathbb{R}$ and any measure-preserving bijection $\sigma$, we have $\mathbb{E}_{\mathsf{M}}f\overset{d}{=}\mathbb{E}_{\mathsf{M}}f\circ(\sigma,\sigma)$. By [26, Prop. 9.1], it further suffices to prove this equality for $\sigma\in\mathsf{S}_{m}$, a permutation of the $m$ intervals $I_{1}^{(m)},\ldots,I_{m}^{(m)}$ acting by $\sigma((i-1)/m+x)=(\sigma(i)-1)/m+x$ for $x\in[0,1/m)$ for $i\in[m]$.
@@ 6 Conclusions and Future Directions
We have introduced a new notion of limits of growing graphs based on convergence of their fixed-size random quotients. We showed that this notion of convergence can be equivalently characterized in terms of convergence of graph parameters such as homomorphism numbers and quotient densities. We characterized limits of quotient-convergent graph sequences via grapheurs, which are certain random exchangeable measures on $0,1]^{2}$. The limiting grapheur of a quotient-convergent sequence of graphs describes the asymptotic distribution of edge weight in the graphs. We then presented another equivalent view of quotient convergence via edge sampling, allowing us to prove an edge-based analog of Szemerédi’s regularity lemma and to test properties of arbitrarily-large graphs by sampling edges from them. The number of edges needed is indpendent of the size of the graph, only depending on the desired accuracy and the Lipschitz constant of the graph parameter in question. Finally, we studied equipartition-consistent random graph models, showing that any such model corresponds to the sequence of random quotients associated to a mixture of grapheurs. We conclude by mentioning several questions suggested by our work.